OpenStudy (anonymous):

Solve: square root x+4=-10

OpenStudy (anonymous):

\[\sqrt{x+4}=-10\]

OpenStudy (matt101):

\[\sqrt{x+4}=-10\]\[x+4=100\]\[x=96\]